Tashin Baker is planning an expansion of a square flower garden in a city park. If each side of the original square is increased by 7 meters and the new total area of the garden will be 144 square meters. Find the lengh of the original garden.
Answer by nerdybill(7384)

Tashin Baker is planning an expansion of a square flower garden in a city park. If each side of the original square is increased by 7 meters and the new total area of the garden will be 144 square meters. Find the lengh of the original garden.
.
Let x = length of original garden
then
(x+7)(x+7) = 144
x^2 + 14x + 49 = 144
x^2 + 14x - 95 = 0
.
Factoring the left side:
(x+19)(x-5) = 0
x = {-19, 5}
We can toss out the negative solution leaving:
x = 5 meters (measure of one side of original garden)